# An Uncertainty Quantification Framework for Deep Learning-Based Automatic Modulation Classification

### v1

@@ III-D Out-of-distribution Samples
To further assess the robustness of our framework and simulate real-world environmental factors, we generate adversarial examples, denoted as $\textbf{r}^{\prime}$, which are considered OOD samples [39]. Deep learning AMC classifiers are highly susceptible to adversarial examples as they induce erroneous predictions, reflected in the softmax output of standalone models, particularly on high SNR signals. Although approaches to mitigate such samples have been investigated [40, 41, 42], they specifically target improving the robustness to adversarial attacks and do not consider the general UQ capabilities of the model. These samples degrade classification performance by pushing the received signal across the decision boundaries of DL models. This phenomenon is often represented as an adversarial perturbation, $\boldsymbol{\delta}\in\mathbb{R}^{\ell\times 2}$, which is introduced to existing samples and formulated as
@@
[…]
@@
In practice it is difficult to analytically solve (8) due to to its excessive non-linearity. Thus, solution to (8) are approximated in practice. In this work, we utilize the Fast Gradient Sign Method (FGSM) [43], to approximate a solution to (8), which is given by
@@
where $\epsilon=||\boldsymbol{\delta}||_{\infty}$ represents the $l_{\infty}$-bounded perturbation magnitude applied to the input samples.
We quantify the effect of the additive perturbation using the perturbation-to-noise ratio (PNR) [24]. Here, we compute the expected value of the power of the perturbation ${\mathbb{E}\left[\|\boldsymbol{\delta}\|_{2}^{2}\right]}$ and then obtain the expected value of the power of the received signal ${\mathbb{E}\left[\|\textbf{r}\|_{2}^{2}\right]}$. Using these quantities, along with the SNR of the received signal, the PNR is given by
@@ IV Performance Evaluation
In this section, we assess the effectiveness of our proposed UQ framework for AMC. We begin by detailing our experimental setup in Sec. IV-A. In Sec. IV-B, we provide an in-depth analysis of the uncertainty estimation capabilities of our proposed framework compared to the proposed baselines. Finally, in Sec. IV-C, we investigate our framework’s robustness to out-of-distribution (OOD) samples and adversarial perturbations, a critical consideration for deploying AMC systems in real-world, dynamic environments.
@@ IV-A Experimental Setup
We perform our empirical evaluation on two AMC datasets: RadioML2016.10a [44] and RadioML2018.01a [11]. RadioML2016.10a is a synthetic dataset generated using GNU radio while RadioML2018.01a is collected over-the-air on a wireless testbed making it a more difficult dataset but a better representation of real-world AMC data. RadioML2016.10a is comprised of $C=11$ modulation types ranging from $-10$ dB to $20$ dB SNR in increments of $2$ dB. For each SNR, the dataset contains $12,000$ signals, where $\ell=128$ samples in length. In comparison, RadioML2018.01a is an over-the-air dataset that contains $C=24$ modulation types, ranging from $-20$ dB to $18$ dB in increments of $2$ dB, with each signal consisting of $\ell=1024$ samples and consists of $19,661$ signals per SNR. Fig. 1 visualizes selected constellations used in our analysis. In each experiment, we use and 80%/20% split of the signals at each SNR to form $\mathcal{X}_{\text{tr}}$ and $\mathcal{X}_{\text{te}}$. $\mathcal{X}_{\text{tr}}$ is used to train each classifier in the ensemble. After training, we report the UQ characteristics of the ensemble on $\mathcal{X}_{\text{te}}$.
Our framework employs an ensemble of $B=15$ CNNs, selected for their demonstrated effectiveness in AMC tasks [29]. All models share the same architecture, which is shown in Table I and are all independently trained on $\mathcal{X}_{\text{tr}}$ by minimizing (3) using stochastic gradient descent (SGD). Each model in the ensemble is trained using 100 epochs, a batch size of 256, and a learning rate of 0.001. Although we use the CNN architecture shown in Table I, our framework can be extended to incorporate an ensemble of classifiers with any architecture. As we will show, our ensemble of AMC classifiers consistently outperforms a standalone AMC classifier with the same architecture.
@@
Output - Softmax $24$
Figure 2: UQ metrics on the 2016.10a (left column) and RML 2018.01a (right column) datasets. We see that our proposed ensemble model outperforms both considered baselines on each considered metric for both datasets.
Figure 3: Correctly predicted CI widths of our proposed ensemble (left column) and the baseline BNN approach (right column) on RML 2018.01a. Here, we see that our ensemble has wider CI widths compared to the BNN, demonstrating that our proposed ensemble approach can characterize uncertainty to a higher extent in comparison to the BNN.
@@ IV-B Uncertainty Quantification
In this section, we examine our framework’s UQ capabilities across a range of SNRs. In addition, we compare the performance of our framework to the two most common approaches for data-driven AMC with UQ: a standalone CNN classifier [10, 7, 9] and an AMC-based BNN [34]. The CNN is a single deep learning classifier that provides state-of-the-art performance but limited UQ capabilities. BNNs are another deep learning-based AMC framework capable of quantifying uncertainty better than standalone deep learning models by generating predictive distributions over point estimates, similar to our proposed ensemble, during inference. Yet, as we will show in our empirical analysis, BNNs suffer in baseline classification performance in comparison to state-of-the-art standalone deep learning classifiers. Contrary to both of these methods, we will see that our proposed deep ensemble framework for AMC provides both state-of-the-art classification performance as well as robust UQ.
@@
Fig. 2 shows the performance of our approach in comparison to each considered baselines in terms of our considered scoring metrics. As shown in Fig. 2, our ensemble-based approach consistently outperforms both a standalone CNN and a BNN across the two datasets. Notably, while the single CNN and the ensemble exhibit relatively stable accuracy in both datasets, the BNN’s accuracy declines sharply on the RML 2018.01a dataset – likely due to the more complex real-world signals in it compared to the GNU radio generated data in the RML 2016.10a dataset —highlighting its limitations in AMC applications. Despite these challenges, the ensemble maintains the highest overall accuracy. Similarly, Fig. 2 shows that our proposed also outperforms the standalone CNN and the BNN, in terms of NLL, consistently achieving a lower NLL score across the entire SNR range. The BNN, in particular, consistently attains a high NLL, relative to our proposed ensemble, in both considered datasets at each considered SNR. Lastly, we see from Fig. 2 that our ensemble consistently achieves the lowest Brier scores in all considered environments, further demonstrating its robustness relative to the other considered baselines.
@@
In Fig. 3, we show the CI widths of low, medium, and high SNR values of our proposed ensemble model in comparison to BNNs on correctly predicted signals. In this scenario, we omit results from the standalone CNN, as they produce point predictions, resulting in CI widths of zero. From Fig. 3, we observe that, at $-10$ dB (low SNR), the BNN exhibits a mix of narrow and moderate CI widths, suggesting some level of uncertainty in low-SNR conditions. However, as the SNR increases to 00 dB and $8$ dB, the BNN’s CI widths shift toward smaller values, indicating greater confidence in its predictions. This behavior aligns with expectations, as higher SNR leads to cleaner input data, reducing uncertainty. In contrast, our proposed ensemble model maintains a higher degree of uncertainty across all SNR levels, with wider CI widths than the BNN. However, as SNR increases, the ensemble also becomes more confident in its correct predictions, while still preserving some uncertainty in certain cases. Notably, our proposed ensemble achieves higher classification accuracy than the BNN (as shown in Fig. 2), despite its broader confidence intervals. This suggests that while the ensemble model expresses more uncertainty, it does not come at the cost of accuracy. Instead, it offers a more calibrated representation of confidence, ensuring that even in high-SNR conditions, some uncertainty is retained where appropriate.
@@
Figure 4: Incorrectly predicted CI widths of our proposed ensemble (left column) and the baseline BNN approach (right column) on RML 2018.01a. Similar to Fig. 3, we see here that our ensemble produces wider CI widths compared to the BNN, demonstrating that our proposed ensemble can more effectively characterize uncertainty estimates, particularly on incorrectly predicted signals.
@@
We now turn our focus to the CI widths of low, medium, and high SNR values of our proposed ensemble model, in comparison to BNNs, on incorrectly predicted signals. In this scenario, robust uncertainty is demonstrated with high CI widths as we would not want a model to express high confidence in a wrong AMC prediction. As shown in Fig.  4, at $-10$ dB, the BNN primarily produces near-zero CI widths, meaning it is overconfident even when it is wrong. This misplaced confidence persists at higher SNRs (e.g., 00 dB and $8$ dB), indicating that the BNN fails to appropriately capture uncertainty in its misclassifications. In contrast, our proposed ensemble exhibits wider CI widths for incorrect predictions, suggesting that it remains more cautious, especially in low-SNR environments, where the performance of deep learning-based AMC classifiers is known to struggle. As the SNR increases, our proposed ensemble’s average CI width remains wide, reflecting an appropriate level of uncertainty even in moderate to low noise conditions. Meanwhile, the BNN’s overconfidence remains evident across all SNR levels, failing to distinguish between correct and incorrect predictions in terms of uncertainty representation.
@@
Next, we assess coverage proportions in Fig. 5, which describe how well each model’s predictive intervals capture the true class. A higher coverage indicates stronger uncertainty representation. Here, we omit the coverage of the single model since it produces point estimates, which results in coverage proportions 00. In Fig. 5, we see that the ensemble outperforms both considered baselines, particularly surpassing the coverage proportions of BNNs. Moreover, we see this trend hold for both the strict and relaxed definitions of coverage (as defined in Sec. III-B), indicating that our proposed ensemble can represent, to a higher degree than the BNN, both its uncertainty associated with its predicted class as well its uncertainty associated with its prediction that the input does not belong to any other class.
@@
Lastly, we evaluate high-confidence sets in Fig. 6, where we label a prediction as high-confidence if the model assigns at least an 80% probability to a particular class (as elaborated on in Sec. III-B). The standalone CNN leads in this category, reflecting its tendency to be overconfident. Conversely, the BNN struggles with underconfidence, consistently producing fewer high-confidence predictions. Our proposed ensemble model strikes a balanced approach, achieving strong accuracy while also retaining suitable uncertainty, even at higher SNR levels for its high-confidence prediction, reinforcing its ability to characterize uncertainty to a higher degree in comparison to the considered baselines.
@@
Figure 6: High confidence proportion of our proposed ensemble in comparison to each considered baseline. In both RML 2016.10a (left) and RML 2018.01a (right), our ensemble maintains a middle level between the CNN and BNN, indicating that the ensemble is neither overconfident nor underconfident compared to the other models.
@@ IV-C Out-of-Distribution Performance
Here, we examine how our framework handles OOD scenarios by examining its effects on adversarial examples. Although they are typically below the noise floor of wireless signals, adversarial perturbations can significantly degrade the performance of neural network–based AMC methods by inducing models to output incorrect predictions due to their underlying shifted distributions [45].
@@
We first examine the effect of adding a constant PNR of 5 dB across the entire SNR range. Fig. 7 displays each model’s accuracy on unperturbed samples for comparison. While the ensemble and the standalone CNN show similar performance on clean signals, the ensemble significantly outperforms the CNN on OOD AMC signals. This disparity highlights the ensemble’s superior UQ capabilities in challenging signal environments. We next examine our proposed ensemble’s accuracy across increasing PNRs at a constant SNR of $10$ dB. As the PNR increases, the perturbed sample is shifted further from its unperturbed counterpart, thereby increasing the chance of misclassification and further shifting the distribution of the adversarial examples. Fig. 8 includes accuracies under normal (unperturbed) conditions, shown as straight lines since SNR remains constant. Here, we see that even as the PNR increases, our proposed ensemble is able to withstand the performance degradation to a greater extent than both the CNN and the BNN. Overall, these findings highlight our proposed ensemble’s ability to retain robust performance under adversarial attacks, reinforcing its ability to characterize UQ to a greater extent in comparison to the considered baselines while simultaneously striking a balance between high confidence and well-calibrated uncertainty.
@@
Figure 7: Accuracy after applying a constant perturbation of $5$ dB across varying SNRs on 2016.10a (left) and 2018.01a (right). The results demonstrate that, under both normal and adversarial conditions, our ensemble model achieves the best performance at high and low SNR values.
@@
Figure 8: Accuracy after varying the perturbation on $10$ dB SNR signals on the 2016.10a (left) and 2018.01a (right) dataset. Here, we see that our proposed ensemble is able to maintain higher performance in comparison to the considered baselines as the PNR increases and approaches the noise floor.

### v2

@@ III-E Out-of-distribution Samples
To further assess the robustness of our framework and simulate real-world environmental factors, we generate adversarial examples, denoted as $\textbf{r}^{\prime}$, which are considered OOD samples [ref56]. Deep learning AMC classifiers are highly susceptible to adversarial examples as they induce erroneous predictions, reflected in the softmax output of standalone models, particularly on high SNR signals. Although approaches to mitigate such samples have been investigated [ade, amc_adv_trn, amc_def2], they specifically target improving the robustness to adversarial attacks and do not consider the general UQ capabilities of the model. These samples degrade classification performance by pushing the received signal across the decision boundaries of DL models. This phenomenon is often represented as an adversarial perturbation, $\boldsymbol{\delta}\in\mathbb{R}^{\ell\times 2}$, which is introduced to existing samples and formulated as
@@
$\text{s.t.}\quad f_{\theta_{b}}(\mathbf{r})\neq f_{\theta_{b}}(\mathbf{r}+\boldsymbol{\delta})$
@@
In practice it is difficult to analytically solve (11) due to its excessive non-linearity. Thus, solution to (11) are approximated in practice. In this work, we utilize the Fast Gradient Sign Method (FGSM) [ref51], to approximate a solution to (11), which is given by
@@
We quantify the effect of the additive perturbation using the perturbation-to-noise ratio (PNR) [amc_adv_atk]. Here, we compute the expected value of the power of the perturbation $\mathbb{E}\left[\|\boldsymbol{\delta}\|_{2}^{2}\right]$} and then obtain the expected value of the power of the received signal $\mathbb{E}\left[\|\textbf{r}\|_{2}^{2}\right]$}. Using these quantities, along with the SNR of the received signal, the PNR is given by
$\text{PNR [dB]}=\frac{\mathbb{E}\left[\|\boldsymbol{\delta}\|_{2}^{2}\right]}{\mathbb{E}\left[\|\textbf{r}\|_{2}^{2}\right]}\,\text{[dB]}+\text{SNR [dB]},$ (13)
@@ IV Performance Evaluation
In this section, we assess the effectiveness of our proposed UQ framework for AMC. We begin by detailing our experimental setup in Sec. IV-A. In Sec. IV-B, we provide an in-depth analysis of the uncertainty estimation capabilities of our proposed framework compared to the proposed baselines. In Sec. IV-C, we investigate our framework’s robustness to OOD samples and adversarial perturbations, a critical consideration for deploying AMC systems in real-world, dynamic environments. Finally, in Sec. IV-D we provide a discussion as well as insights highlighting the benefits of our proposed method.
@@ IV-A Experimental Setup
We perform our empirical evaluation on two AMC datasets: RadioML2016.10a [ref31] and RadioML2018.01a [ref32]. RadioML2016.10a is a synthetic dataset generated using GNU radio while RadioML2018.01a is collected over-the-air on a wireless testbed making it a more difficult dataset but a better representation of real-world AMC data. RadioML2016.10a is comprised of $C=11$ modulation types ranging from 10-10 dB to $20$ dB SNR in increments of $2$ dB. For each SNR, the dataset contains $12,000$ signals, where $\ell=128$ samples in length. In comparison, RadioML2018.01a is an over-the-air dataset that contains $C=24$ modulation types, ranging from 20-20 dB to $18$ dB in increments of $2$ dB, with each signal consisting of $\ell=1024$ samples and consists of $19,661$ signals per SNR. Fig. 1 visualizes selected constellations used in our analysis. In each experiment, we use an 80%/20% split of the signals at each SNR to form $\mathcal{X}_{\text{tr}}$ and $\mathcal{X}_{\text{te}}$. $\mathcal{X}_{\text{tr}}$ is used to train each classifier in the ensemble. After training, we report the UQ characteristics of the ensemble on $\mathcal{X}_{\text{te}}$.
Our framework employs an ensemble of $B=15$ CNNs, selected for their demonstrated effectiveness in AMC tasks [ref9]. All models share the same architecture, which is shown in Table I and are all independently trained on $\mathcal{X}_{\text{tr}}$ by minimizing (3) using stochastic gradient descent (SGD). Each model in the ensemble is trained for 100 epochs with a batch size of 256 and a learning rate of 0.001. The hyperparameters were initially selected based on commonly used settings in the AMC literature [ref_bs], and they were further fine-tuned for our application using a grid search. Although we use the CNN architecture shown in Table I, our framework can be extended to incorporate an ensemble of classifiers with any architecture. As we will show, our ensemble of AMC classifiers consistently outperforms a standalone AMC classifier with the same architecture.
@@
Output - Softmax 2424
Figure 2: UQ metrics on the 2016.10a (left column) and RML 2018.01a (right column) datasets. We see that our proposed ensemble model outperforms all baselines on each considered metric for both datasets.
Figure 3: Correctly predicted CI widths of our proposed ensemble (left column), the baseline BNN approach (middle column), and the baseline weighted ensemble (right column) on RML 2018.01a. Here, we see that our ensemble has wider CI widths compared to both baselines, demonstrating that our proposed ensemble approach can characterize uncertainty to a higher extent in comparison to the BNN and weighted ensemble.
@@ IV-B Uncertainty Quantification
In this section, we examine our framework’s UQ capabilities across a range of SNRs. In addition, we compare the performance of our framework to the two most common approaches for data-driven AMC with UQ: a standalone CNN classifier [ref29, ref33, ref35] and an AMC-based BNN [ref2]. The CNN is a single deep learning classifier that provides state-of-the-art performance but limited UQ capabilities. BNNs are another deep learning-based AMC framework capable of quantifying uncertainty better than standalone deep learning models by generating predictive distributions over point estimates, similar to our proposed ensemble, during inference. Yet, as we will show in our empirical analysis, BNNs suffer in baseline classification performance in comparison to state-of-the-art standalone deep learning classifiers. Furthermore, we introduce an extra baseline: An SNR-aware weighted ensemble inspired by [ref_weighted], where we use an SNR-aware approach to train each model in an ensemble on a specific SNR and then calculate the weight of each model using the Shannon entropy. This approach introduces a baseline in which each model in the ensemble is trained on a unique SNR, thus distinguishing between epistemic uncertainty (model uncertainty) and aleatoric uncertainty (channel noise) in AMC tasks. Here, models with high entropy (e.g., classifiers trained on low SNR signals) are given a lower weight and models with low entropy (e.g., classifiers trained on high SNR signals) are given more weight (proportional to the Shannon entropy of each classifier). Note that although [ref_weighted] does not explicitly apply a weighted ensemble to AMC, it shows the potential of a weighted ensemble to characterize aleatoric uncertainty so we adapt it for AMC as an additional baseline. Contrary to all of these methods, we will see that our proposed deep ensemble framework for AMC provides both state-of-the-art classification performance as well as robust UQ.
@@
Fig. 2 shows the performance of our approach in comparison to each considered baselines in terms of our considered scoring metrics. As shown in Fig. 2, our ensemble-based approach consistently outperforms a standalone CNN, the BNN, and the weighted ensemble on both considered datasets. Notably, while the single CNN, the weighted ensemble, and our proposed ensemble exhibit relatively stable accuracy in both datasets, the BNN’s accuracy declines sharply on the RML 2018.01a dataset – likely due to the more complex real-world signals in it compared to the GNU radio generated data in the RML 2016.10a dataset, highlighting its limitations in AMC applications. Despite these challenges, our proposed ensemble maintains the highest overall accuracy. Similarly, Fig. 2 shows that our proposed ensemble also outperforms the standalone CNN, the BNN, and the weighted ensemble in terms of NLL, consistently achieving a lower NLL score across the entire SNR range. The BNN, in particular, consistently attains a high NLL, relative to our proposed ensemble, in both considered datasets at each considered SNR. Lastly, we see from Fig. 2 that our ensemble consistently achieves the lowest Brier scores in all considered environments, further demonstrating its robustness relative to the other considered baselines.
@@
In Fig. 3, we show the CI widths of low, medium, and high SNR values of our proposed ensemble model in comparison to BNNs and the weighted ensemble on correctly predicted signals. In this scenario, we omit results from the standalone CNN, as they produce point predictions, resulting in CI widths of zero. From Fig. 3, we observe that, at 10-10 dB (low SNR), the BNN exhibits a mix of narrow and moderate CI widths, suggesting some level of uncertainty in low-SNR conditions. However, as the SNR increases to 0 dB and $8$ dB, the BNN’s CI widths shift toward smaller values, indicating greater confidence in its predictions. This behavior aligns with expectations, as higher SNR leads to cleaner input data, reducing uncertainty. The weighted ensemble, on the other hand, attains narrower widths at low SNR, where higher uncertainty expression is more crucial, and higher widths at high SNR, where higher uncertainty expression is less crucial. In contrast, our proposed ensemble model maintains a higher degree of uncertainty across all SNR levels, with wider CI widths than both considered baselines. Moreover, as SNR increases, the ensemble also becomes more confident in its correct predictions, while still preserving some uncertainty in certain cases. Notably, our proposed ensemble achieves higher classification accuracy than the BNN and the weighted ensemble baseline (as shown in Fig. 2), despite its broader confidence intervals. This suggests that while the ensemble model expresses more uncertainty, it does not come at the cost of accuracy. Instead, it offers a more calibrated representation of confidence, ensuring that even in high-SNR conditions, some uncertainty is retained where appropriate.
@@
Figure 4: Incorrectly predicted CI widths of our proposed ensemble (left column), the baseline BNN approach (middle column), and the baseline weighted ensemble (right column) on RML 2018.01a. Similar to Fig. 3, we see here that our ensemble produces wider CI widths compared to both baselines, demonstrating that our proposed ensemble can more effectively characterize uncertainty estimates, particularly on incorrectly predicted signals.
@@
We now turn our focus to the CI widths of low, medium, and high SNR values of our proposed ensemble model on incorrectly predicted signals. In this scenario, robust uncertainty is demonstrated with high CI widths as we would not want a model to express high confidence in a wrong AMC prediction. As shown in Fig.  4, at 10-10 dB, the BNN and the weighted ensemble baseline primarily produce near-zero CI widths, meaning they are overconfident even when they are incorrect. This misplaced confidence persists at higher SNRs (e.g., 0 dB and $8$ dB), indicating that both baselines fail to appropriately capture uncertainty in their misclassifications. In contrast, our proposed ensemble exhibits wider CI widths for incorrect predictions, suggesting that it remains more cautious, especially in low-SNR environments, where the performance of deep learning-based AMC classifiers is known to struggle. As the SNR increases, our proposed ensemble’s average CI width remains wide, reflecting an appropriate level of uncertainty even in moderate to low noise conditions. Meanwhile, the overconfidence of the other baselines remains evident across all SNR levels, failing to distinguish between correct and incorrect predictions in terms of uncertainty representation.
@@
Next, we assess coverage proportions in Fig. 5, which describe how well each model’s predictive intervals capture the true class. A higher coverage indicates stronger uncertainty representation. Here, we omit the coverage of the single model since it produces point estimates, which results in coverage proportions of 0. In Fig. 5, we see that the ensemble outperforms both considered baselines, particularly surpassing the coverage proportions of BNNs. Moreover, we see this trend hold for both the strict and relaxed definitions of coverage (as defined in Sec. III-D), indicating that our proposed ensemble can represent, to a higher degree than each considered baseline, both its uncertainty associated with its predicted class as well as its uncertainty associated with its prediction that the input does not belong to any other class.
@@
We now evaluate the high-confidence sets in Fig. 6, where we label a prediction as high-confidence if the model assigns at least an 80% probability to a particular class (as elaborated on in Sec. III-D). The standalone CNN leads in this category, reflecting its tendency to be overconfident. Conversely, both the BNN and the weighted ensemble baselines struggle with underconfidence, consistently producing fewer high-confidence predictions. Our proposed ensemble model strikes a balanced approach, achieving strong accuracy while also retaining suitable uncertainty, even at higher SNR levels for its high-confidence prediction, reinforcing its ability to characterize uncertainty to a higher degree in comparison to the considered baselines.
Finally, we analyze the UQ capabilities of our framework by measuring the ECE and KL Divergence. A lower ECE indicates better model calibration, and as shown in Fig. 7, our ensemble consistently achieves the smallest calibration error across all SNR levels, whereas the BNN spikes and trails behind. Similarly, the weighted ensemble baseline exhibits a higher degree of ECE across the entire SNR range, indicating its confidence estimates are not well-calibrated to its accuracy. The same trend holds for the KL divergence, where our ensemble maintains the lowest divergence from the true predictive distribution, outperforming both the standalone CNN, the BNN, and the weighted ensemble baselines showcasing more expressive and robust UQ capabilities.
In the context of AMC, it is useful to distinguish between epistemic and aleatoric uncertainty. Epistemic uncertainty arises from model limitations (e.g., insufficient training data) and can potentially be reduced by larger training sets or higher-parameter models. Aleatoric uncertainty, on the other hand, stems from inherent noise in the received signals (e.g., low SNR or OOD data) and is a result of the quality of the received signals used for training. Our results shown in Figs. 27 demonstrate that our proposed ensemble-based framework captures both types of uncertainty: the diversity among independently trained models reflects epistemic uncertainty, while the variability in predictions across signal conditions (e.g., at low SNRs) captures aleatoric uncertainty. Compared to the BNN and weighted ensemble baselines, our method more effectively disentangles these sources, as shown by its higher confidence interval widths on misclassified samples (epistemic) and its improved calibration under noisy conditions (aleatoric).
@@
Figure 6: High confidence proportion of our proposed ensemble in comparison to each considered baseline. In both RML 2016.10a (left) and RML 2018.01a (right), our ensemble maintains a middle level between the CNN, baseline BNN, and baseline weighted ensemble indicating that our proposed ensemble is neither overconfident nor underconfident and able to strike a UQ balance compared to the other models.
Figure 7: ECE and KL Divergence of our proposed ensemble in comparison to each considered baseline on the 2016.10a (left) and 2018.01a (right) dataset. Here, we see that our proposed ensemble is able to achieve the lowest ECE and KL divergence compared to the baseline showing improved calibration.
@@ IV-C Out-of-Distribution Performance
Here, we examine how our framework handles OOD scenarios by examining its effects on adversarial examples. Although they are typically below the noise floor of wireless signals, adversarial perturbations can significantly degrade the performance of neural network–based AMC methods by inducing models to output incorrect predictions due to their underlying shifted distributions [ref54].
@@
We first examine the effect of adding a constant PNR of 5 dB across the entire SNR range. Fig. 8 displays each model’s accuracy on unperturbed samples for comparison. While the ensemble and the standalone CNN show similar performance on clean signals, the ensemble significantly outperforms the CNN on OOD AMC signals. This disparity highlights the ensemble’s superior UQ capabilities in challenging signal environments. We next examine our proposed ensemble’s accuracy across increasing PNR at a constant SNR of $10$ dB. As the PNR increases, the perturbed sample is shifted further from its unperturbed counterpart, thereby increasing the chance of misclassification and further shifting the distribution of the adversarial examples. Fig. 9 includes accuracies under normal (unperturbed) conditions, shown as straight lines since SNR remains constant. Here, we see that even as the PNR increases, our proposed ensemble is able to withstand the performance degradation to a greater extent than both the CNN and the BNN. Overall, these findings highlight our proposed ensemble’s ability to retain robust performance under adversarial attacks, reinforcing its ability to characterize UQ to a greater extent in comparison to the considered baselines while simultaneously striking a balance between high confidence and well-calibrated uncertainty.
@@
Figure 8: Accuracy after applying a constant perturbation of 55 dB across varying SNRs on 2016.10a (left) and 2018.01a (right). The results demonstrate that, under both normal and adversarial conditions, our ensemble model achieves the best performance at high and low SNR values.
@@
Figure 9: Accuracy after varying the perturbation on 1010 dB SNR signals on the 2016.10a (left) and 2018.01a (right) dataset. Here, we see that our proposed ensemble is able to maintain higher performance in comparison to the considered baselines as the PNR increases and approaches the noise floor.
IV-D Discussion
Here, we examine why our proposed ensemble outperforms BNNs, which are specifically designed to deliver robust UQ. While BNNs theoretically offer a Bayesian measure of uncertainty via their posterior distribution-based estimates, they often require unrealistic approximations that limit their ability to attain a high baseline accuracy in AMC tasks. For example, BNNs use variational inference, which by definition assumes independence between weights in the model. However, in AMC, the model operates on raw time-domain IQ samples, which are highly correlated with their nearby features. As a result, BNNs become limited from achieving state-of-the-art accuracy on AMC data as the correlation between features are highly indicative in discerning modulation type and BNNs are unable to exploit such correlations for high accuracy. In contrast, CNNs are not limited by the assumption of feature independence and, thus, ensembles of independently trained CNNs are able to exploit spatial regions of the received waveform (i.e., various portions of the time-domain signal) and are able to learn distinctive features between signals, resulting in robust uncertainty estimates through implicit posterior approximations. This makes our proposed ensemble particularly effective in comparison to BNNs, despite lacking a fully Bayesian foundation.
Furthermore, BNNs assume Gaussian priors on their weights (often with standard zero-mean, unit-variance distributions), which tends to break down for AMC as each feature in a received waveform is not necessarily Gaussian in real-world radio data, and moreover, AMC data may require more accurate assumptions on their priors to achieve high accuracy. For example, the channel distribution of received signals could differ from priors, hindering the ability of BNNs to effectively learn distinguishing features if an incorrect channel distribution assumption is made. CNNs, in contrast require no such prior assumption on the channel distribution of received signals and are not limited to learning weights of a certain distribution, allowing more expressive power and directly allowing higher AMC performance without specific channel assumptions. Thus, ensembles of CNNs, as proposed in our framework, not only achieve higher classification performance compared to BNNs, but they also express uncertainty more effectively due to their stronger expressive abilities.
Finally, as shown in Fig. 2 – Fig. 9, our ensemble consistently outperforms the BNN baseline in both predictive accuracy and uncertainty calibration metrics (e.g., NLL and ECE). These results support the hypothesis that our proposed deep ensemble-based UQ approach offers more robust performance, even if BNNs offer a theoretically principled approach to UQ.